OpenStudy (anonymous):

please help! A cube shaped box has a volume of 512 cubic inches.. Find the length of each side given . v=s^3

OpenStudy (anonymous):

v=s^3 512=s^3 s=(512)^(1/3)

OpenStudy (anonymous):

Solve for s

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

i gace b noo calculator

OpenStudy (anonymous):